# Identifying treatment response subgroups in observational time-to-event data

### v4

@@ Acknowledgements
Appendix A Proof
A.1 Individualised Treatment Effect
This section derives the individualised treatment effect expression introduced in Section 3.
@@ A.1 Individualised Treatment Effect
$=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)\mid X=x)$
$-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
$=\mathbb{P}($ $T^{\prime}\geq t\mid A=1,X=x)$
$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under assumption 1 and 3)
Appendix B Synthetic analysis
B.1 Data generation
We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to (Nagpal et al., 2022a). The following describes our generation process:
@@ B.1 Data generation
Each patient’s membership $Z$ is drawn from a multinomial with equal probability. Group membership informs the two first covariates through the parametrisation of the bivariate normal distribution with centres $c_{k}$ equal to $(0,2.25)$, $(-2.25,-1)$, and $(2.25,-1)$. All other covariates are drawn from standard normal distributions. Formally, this procedure is described as:
@@
with MVN denoting a multivariate normal distribution, $c^{\prime}_{k}$ random cluster centres, and $I^{n}$, the identity covariance matrix of dimension $n$. Note that we introduce $Y$ to present a covariates structure that is independent of the treatment responses.
(a) $K=2$
(b) $K=3$
(c) $K=4$
Figure 5: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross validated average estimated treatment effect. Lines in grey corresponds to the ground truth.
@@
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world studies. Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as Randomised and Observational. Randomised consists of a Bernoulli draw using the realisation of $P$. Observational reflects an assignment dependent upon the observed covariates.
@@
B.2 Training and hyperparameter optimisation
@@
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search. All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
@@
B.3 Sensitivity to choice of K
In Section 4, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$. While this outcome-guided selection of the number of clusters is a core strength of the proposed methodology in comparison to existing strategies, we explore how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters. This analysis examines the risk associated with misspecifying the number of clusters.
@@ B.3 Sensitivity to choice of K
Table 2: Cross-validated performance (with standard deviation in parenthesis).
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.643 (0.067) 0.060 (0.016) 0.021 (0.007) 0.070 (0.013) CSC Unadjusted 0.643 (0.071) 0.064 (0.013) 0.018 (0.008) 0.060 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.788 (0.113) 0.040 (0.017) 0.041 (0.017) 0.086 (0.043) CSC Unadjusted 0.626 (0.170) 0.080 (0.015) 0.031 (0.009) 0.095 (0.049) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
Table 3: Cross-validated performance (with standard deviation in parenthesis) when clusters are of different sizes.
Figure 5 illustrates the cross-validated subgroup treatment effects when the model is trained with $K=2,3$ and $4$ clusters, despite the existence of 3 clusters. When using a number of clusters different than indicated by the elbow heuristic, the identified clusters are less stable as not capturing the true underlying treatment responses, as shown by the increased standard deviation. Uncertainty increases with misspecified models, providing an additional tool to select the appropriate number of cluster: the stability of the estimated clusters over cross-validation indicates a better alignment with underlying distribution.
B.4 Linear Modelling
While the flexibility of neural network allows flexibility of the survival distribution, one could consider more interpretable assignment $G$ and treatment $W$ components than the neural network approach proposed in the main text. This section explores the recovery of treatment effects and subgroups when using linear regressions, i.e. no hidden layers in a neural network, instead of the previous neural networks.
Table 2 summarises the results presented in the main paper with the additional linear alternative in which both the treatment and assignment networks are replaced by a linear model, referred as CSC Linear. These results demonstrate that enforcing linear relation between covariates and the different quantities of interest can reduce the risk of overfitting in small datasets as shown by the improved Rand-Index. Note that under larger dataset or more complex relation between covariates, group membership and treatment, one should consider more flexible modelling, such as the proposed neural approach.
B.5 Alternative data generations
@@
B.5.1 Unequal cluster size
Section B.1 presents an equally-distributed population over the different clusters. In medical settings, the underlying subgroups may differ in size. This section explores an alternative scenario in which the population is distributed over the three clusters as follows: 62.5%, 25% and 12.5%.
@@ B.5.1 Unequal cluster size
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.310 (0.187) 0.236 (0.103) 0.110 (0.033) 0.149 (0.106) CSC Unadjusted 0.443 (0.285) 0.176 (0.098) 0.087 (0.046) 0.132 (0.112) CMHE ($K=3$) 0.113 (0.100) 0.296 (0.040) 0.137 (0.042) 0.188 (0.080) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.079 (0.119) 0.359 (0.068) 0.182 (0.089) 0.240 (0.058) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.764 (0.015) 0.027 (0.008) 0.011 (0.001) 0.047 (0.008) CSC Unadjusted 0.729 (0.022) 0.033 (0.008) 0.010 (0.003) 0.039 (0.009) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
Table 4: Cross-validated performance (with standard deviation in parenthesis) with varying $N$ under observational treatment setting.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.830 (0.047) 0.051 (0.008) 0.023 (0.013) 0.020 (0.005) CSC Unadjusted 0.857 (0.033) 0.045 (0.006) 0.030 (0.015) 0.015 (0.003) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.134 (0.090) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.718 (0.198) 0.061 (0.050) 0.057 (0.048) 0.023 (0.005) CSC Unadjusted 0.804 (0.045) 0.037 (0.006) 0.023 (0.008) 0.025 (0.007) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.318 (0.140) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012)
Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates under observational settings.
B.5.2 Number of points
@@
Section B.1 describes a data generation with $3,000$ patients. This section presents results when there are $300$ and $30,000$ patients. A smaller number of patients may impact the neural network capacity to identify underlying subgroups of treatment effect. A larger number makes non-randomisation less of a concern.
Table 4 presents the performance with these different population sizes under an observational treatment assignment. A first observation is that all methodologies present better performance with larger number of points. Further, baselines that do not account for the assignment mechanisms present lower performance with $N=300$ as non-randomisation has an increased impact on treatment effect estimate with a smaller population. This difference decreases when $N=30,000$ with the virtual twins approach presenting the best recovery of the treatment effects. However, throughout the different settings, the proposed CSC presents the best Rand-Index indicating a good recovery of the underlying structure.
B.5.3 Impact of treatment rate
Section B.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.
@@ B.5.3 Impact of treatment rate
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.559 (0.114) 0.024 (0.009) 0.030 (0.017) 0.030 (0.006) CSC Unadjusted 0.481 (0.069) 0.026 (0.005) 0.041 (0.026) 0.032 (0.009) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.254 (0.138) 0.074 (0.027) 0.052 (0.021) 0.149 (0.018) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.584 (0.151) 0.015 (0.007) 0.028 (0.024) 0.035 (0.003) CSC Unadjusted 0.556 (0.249) 0.023 (0.008) 0.051 (0.025) 0.040 (0.006) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.293 (0.125) 0.063 (0.012) 0.051 (0.017) 0.140 (0.011) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
Table 6: Cross-validated performance (with standard deviation in parenthesis) when all covariates clusters presents different treatment responses. This setting aligns with the assumptions made by the step-wise approaches.

### v5

@@ Acknowledgements
Appendix A Related work
While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a, b; Jeanselme et al., 2022) to identify different phenotypes of patients, the ML literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current focus is on estimating population or individual treatment effects in the static (Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020) and survival settings (Katzman et al., 2018; Curth et al., 2021). While these approaches estimate ATE, CATE, or even ITE, they do not directly identify subgroups that may benefit or be harmed by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
Tree-based approaches were proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey and Imbens (2015, 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager and Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect Lipkovich et al. (2011); Loh et al. (2015). Alternatively, McFowland III et al. (2018) propose pattern detection and Wang and Rudin (2022), causal rule set learning to uncover these subgroups. However, all these approaches rely on a local optimisation criterion (Lipkovich and Dmitrienko, 2014) and greedy split exploration. Recently, Nagpal et al. (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models. Our work proposes to address these limitations through a joint optimisation, while avoiding parametric treatment response and proportional hazard assumptions.
Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with survival outcomes that our work explores. At the intersection with survival analysis, Zhang et al. (2017) extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. (2021) propose Bayesian additive models and Zhu and Gallego (2020) propose a step-wise approach with propensity weighting to study observational data. Closest to our work, (An et al., 2023; Jia et al., 2021; Nagpal et al., 2023; Perrin et al., 2024) propose to uncover subgroups within RCTs with survival outcomes. Perrin et al. (2024) compares different recursive and tree-based appraoches to perform this task in RCTs. An et al. (2023) introduces an expectation-maximisation approach to jointly fit a logistic model for subgroup attribution and a Cox model for survival estimation. Jia et al. (2021) propose a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework. Similarly, Nagpal et al. (2023) stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational settings.
Appendix B Proof
This section derives the conditional average treatment effect expression introduced in Sec. 2.
@@ Appendix B Proof
$=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)\mid X=x)-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
$=\mathbb{P}($ $T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under Asmp. 1 and 3)
Appendix C Synthetic analysis
C.1 Data generation
We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to Nagpal et al. (2022a). The following describes our generation process:
@@ C.1 Data generation
Each patient’s membership $Z$ is drawn from a multinomial with equal probability. Group membership informs the first two covariates through the parametrisation of the bivariate normal distribution with centres $c_{k}$ equal to $(0,2.25)$, (2.25,1)(-2.25,-1), and (2.25,1)(2.25,-1). All other covariates are drawn from standard normal distributions. Formally, this procedure is described as:
@@
with MVN denoting a multivariate normal distribution, $c^{\prime}_{k}$ random cluster centres, and $I^{n}$, the identity covariance matrix of dimension $n$. Note that we introduce $Y$ to present a covariate structure that is independent of the treatment responses.
[…]
Figure 6: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross-validated average estimated treatment effect. Lines in grey correspond to the ground truth.
@@
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world studies. Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting an RCT and an observational setting, denoted as Randomised and Observational. Randomised consists of a Bernoulli draw using the realisation of $P$. Observational reflects an assignment dependent upon the observed covariates.
@@
C.2 Training and hyperparameter optimisation
@@
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search. All models were optimised for 1000 epochs using an Adam optimiser (Kingma and Ba, 2015) with early stopping.
@@
C.3 Sensitivity to choice of K
In Sec. 3, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$. While this outcome-guided selection of the number of clusters is a core strength of the proposed methodology in comparison to existing strategies, we explore how the identified treatment effects change under a misspecified model that estimates 2 and 4 clusters. This analysis examines the risk associated with misspecifying the number of clusters.
@@ C.3 Sensitivity to choice of K
Table 3: Cross-validated performance (with standard deviation in parentheses).
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.643 (0.067) 0.060 (0.016) 0.021 (0.007) 0.070 (0.013) CSC Unadjusted 0.643 (0.071) 0.064 (0.013) 0.018 (0.008) 0.060 (0.010) CMHE ($L=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($M=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($M=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.788 (0.113) 0.040 (0.017) 0.041 (0.017) 0.086 (0.043) CSC Unadjusted 0.626 (0.170) 0.080 (0.015) 0.031 (0.009) 0.095 (0.049) CMHE ($L=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($M=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($M=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
Table 4: Cross-validated performance (with standard deviation in parentheses) when clusters are of different sizes.
Figure 6 illustrates the cross-validated subgroup treatment effects when the model is trained with $K=2,3$, and $4$ clusters, despite the existence of 3 clusters. When using a number of clusters different from that indicated by the elbow heuristic, the identified clusters are less stable, as they do not capture the true underlying treatment responses, as shown by the increased standard deviation. Uncertainty increases with misspecified models, providing an additional tool to select the appropriate number of clusters: the stability of the estimated clusters over cross-validation indicates a better alignment with the underlying distribution.
C.4 Linear modelling
While the flexibility of neural networks allows flexibility of the survival distribution, one could consider more interpretable assignment $G$ and treatment $W$ components than the neural network approach proposed in the main text. This section explores the recovery of treatment effects and subgroups when using linear regressions, i.e. no hidden layers in a neural network, instead of the previous neural networks.
Table 3 summarises the results presented in the main paper with the additional linear alternative in which both the treatment and assignment networks are replaced by a linear model, referred as CSC Linear. These results demonstrate that enforcing a linear relation between covariates and the different quantities of interest can reduce the risk of overfitting in small datasets, as shown by the improved Rand-Index. Note that under a larger dataset or a more complex relation between covariates, group membership, and treatment, one should consider more flexible modelling, such as the proposed neural approach.
C.5 Alternative data generations
@@
C.5.1 Unequal cluster size
Sec. C.1 presents an equally distributed population over the different clusters. In medical settings, the underlying subgroups may differ in size. This section explores an alternative scenario in which the population is distributed over the three clusters as follows: 62.5%, 25%, and 12.5%.
@@ C.5.1 Unequal cluster size
C.5.2 Number of points
@@
Sec. C.1 describes a data generation with $3,000$ patients. This section presents results for $300$ and $30,000$ patients. A smaller number of patients may impact the neural network capacity to identify underlying subgroups of treatment effect. A larger number makes non-randomisation less of a concern.
@@
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.310 (0.187) 0.236 (0.103) 0.110 (0.033) 0.149 (0.106) CSC Unadjusted 0.443 (0.285) 0.176 (0.098) 0.087 (0.046) 0.132 (0.112) CMHE ($L=3$) 0.113 (0.100) 0.296 (0.040) 0.137 (0.042) 0.188 (0.080) CMHE ($M=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($M=L$) 0.079 (0.119) 0.359 (0.068) 0.182 (0.089) 0.240 (0.058) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.764 (0.015) 0.027 (0.008) 0.011 (0.001) 0.047 (0.008) CSC Unadjusted 0.729 (0.022) 0.033 (0.008) 0.010 (0.003) 0.039 (0.009) CMHE ($L=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($M=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($M=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
Table 5: Cross-validated performance (with standard deviation in parentheses) with varying NN under an observational treatment setting.
Table 5 presents the performance with these different population sizes under an observational treatment assignment. A first observation is that all methodologies present better performance with a larger number of points. Further, baselines that do not account for the assignment mechanisms present lower performance with $N=300$ as non-randomisation has an increased impact on treatment effect estimates with a smaller population. This difference decreases when $N=30,000$ with the virtual twins approach presenting the best recovery of the treatment effects. However, throughout the different settings, the proposed CSC presents the best Rand-Index, indicating a good recovery of the underlying structure.
C.5.3 Impact of treatment rate
Sec. C.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.